OpenStudy (anonymous):

2x-7/x-5 >0 help please?

OpenStudy (anonymous):

what are you trying to find?

OpenStudy (anonymous):

the two points I'm guessing the answer is\[(-\infty,\ -\frac{ 7 }{ 2 }) \cup (5,\infty)\]

OpenStudy (dtan5457):

Multiply the fraction by (x+5), each. You cross out the bottom part and end up with 2x-7(x+5). That's 2x^2-3x-35>0. Factor that.

OpenStudy (dtan5457):

That becomes (x-5)(2x+7).

OpenStudy (anonymous):

thank u !